# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v4

@@ 7.2 Privacy-Preserving Techniques
8 Compliance to AI Governance and Policy in Agentic AI Systems
@@ 8.1 Regulatory Standards
  • NIST AI Risk Management Framework (AI RMF) [171]. A voluntary U.S. framework organized around the Govern, Map, Measure, Manage functions. It emphasizes organizational governance, risk identification and measurement, and operational risk treatments. Mapping to the RMF also benefits from NISTIR 8312 (Explainable AI principles) for transparency obligations.
  • EU AI Act [125]. A comprehensive EU regulation with phased application: bans on certain practices apply after six months (Feb 2025), obligations for general-purpose AI begin Aug 2025, and most high-risk system obligations (e.g., risk management, documentation, human oversight, post-market monitoring) start Aug 2026. High-risk systems require ongoing risk assessment, technical documentation (Annex IV), logging/traceability (Arts. 11–12), data governance (Art. 10), human oversight (Art. 14), and robustness/accuracy (Art. 15).
@@ 8.2 Auditability
Multi-agent systems can produce emergent and opaque behaviors; auditors must reconstruct the chain of decisions.
@@ 8.4 Documentation & Record-Keeping (Evidence)
  • Technical documentation (EU AI Act [1]): system purpose and risk class, training/eval data governance, performance, limitations, foreseeable misuse, human oversight measures.
@@ 8.5 Post-Market Monitoring & Incident Response
  • Periodic reassessment. Re-run impact/risk assessments (ISO/IEC 42005 [129], ISO/IEC 23894 [130] ) when capabilities, data, or context change materially.
@@ 8.6 Supply-Chain & Secure Development (GenAI)
  • OWASP LLM risks. Mitigate prompt injection, insecure output handling, data poisoning, model DoS, and supply-chain risks using the OWASP [137] Top 10 for LLM applications.
@@ 9 Discussion
In this section, we present our findings and outline directions for future work.
@@ 9.1 Technical Implications of TRiSM for Agentic AI Design
Rather than treating LLM agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” [172]. For example, there is discussion on designing specialized “guardian agents” within AMAS [42]. In this paradigm, such agents serve as proactive monitors that filter sensitive data and establish baselines of normal behavior, while operator agents dynamically enforce policies at runtime (e.g. blocking disallowed actions such as outputting personally identifiable information).
Prior research highlights risks like “excessive agency” [138] where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals). TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds. Likewise, emerging threats specific to Agentic AI systems such as prompt injection attacks, memory poisoning , or cascading hallucinations, highlights the need for more built in risk controls. By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g. masking a sensitive datum or stopping an unsafe action).
9.2 Ethical and Societal Ramifications of Multi-Agent AI
Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions. Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness [173], all of which are vital for public trust in AI systems. Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”. This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed . Recent guidance on trustworthy AI [174] often highlights accountability and explainability as key pillars of trust.
In practice, this review sheds some light on how each autonomous agent decisions should be transparent enough to be understood and challenged by human reviewers when necessary [175]. Human oversight is another ethical imperative tightly coupled with accountability. TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under a defined governance framework [176].
In AMAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly . Therisk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [177]. TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision.
@@ 9.3 Alignment with Emerging AI Regulations and Standards
The principles embedded in TRiSM align closely with emerging regulatory frameworks for AI. This convergence means that adopting TRiSM-based governance can help organizations meet new legal obligations and industry standards. For example, the European Union’s AI Act [125] (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems. These are precisely the capabilities that a TRiSM approach promotes.
Deploying Agentic AI systems responsibly requires a multifaceted governance approach. By adhering to established regulatory standards (and keeping abreast of new ones), by ensuring auditability for transparency and accountability, and by enforcing policies through technical and organizational means, we can significantly mitigate the risks posed by autonomous AI agents. Moreover, embracing evolving best practices – from governance boards to adaptive, code-driven policies – will help maintain control over these systems as they grow in complexity. The end goal is to harness the benefits of Agentic AI (increased efficiency, creativity, and problem-solving at scale) while staying within legal, ethical, and safe bounds, thereby preserving public trust and meeting compliance obligations.
9.4 Limitations and Current Research Gaps
While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research. First, the classical AI domains that works on standard test suites, there is no consensus on how to measure an AI Agent vs the Agentic AI ability to operate safely under TRiSM principles. This makes it difficult to compare different governance strategies or to track progress objectively. We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [134] or “red-teaming” exercises [154] that can stress-test Agentic AI systems and score their resilience.
There is also a lack of real-world validation for many TRiSM-inspired controls. Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [65]. It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments. Another critical gap is adversarial learning. As we improve defenses, attackers will inevitably adapt [178]. Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [38].
Beyond technical aspects, the organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures. Many organizations lack the necessary skill sets or frameworks, making TRiSM adoption superficial or inconsistent. Without a strong organizational commitment, even the best technical framework can falter.
@@ 9.5 Future Roadmap for Agentic AI TRiSM
The community should create open benchmarks and challenge environments to test AMAS governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed Agentic AI systemthat performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” [179], and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [180] , can help discover vulnerabilities in a controlled way before real adversaries do.
We encourage designing better interfaces and protocols for human oversight of Agentic AI systems. Borrowing from human-computer interaction [181] and cognitive engineering [182], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations. There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
A cross-cutting concept relevant to both Governance and Lifecycle Management, controllability captures the ability of humans and oversight systems to steer or override agentic AI. Mechanisms include stop/pause functions, rollback gates role-based oversight, and auditability of inter-agent actions. In practice, controllability is critical to ensuring accountability and preventing runaway behaviors in multi-agent workflows.
@@ 10 Conclusion
The rapid advancement of Agentic AI, driven by AMAS, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services. However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations. This review has explored these challenges, proposing a TRiSM framework structured around five pillars: Explainability, ModelOps, Security, Privacy and their lifecycle Governance, adapted for LLM-based AMAS. Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency. The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities. Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination. By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.
Acknowledgements
Funding for the research was provided through Horizon Europe project AIXpert: An agentic, multi-layer, GenAI-powered backbone to make an AI system explainable, accountable, and transparent (ID:101214389).

### v5

@@ 6.2 Privacy-Preserving Techniques
6.3 Evaluation Benchmarks and Metrics for Agentic AI Security
Evaluating the security and robustness of AMAS requires standardized benchmarks and well-defined metrics to assess both attack vulnerabilities and defense effectiveness. This subsection reviews current evaluation frameworks and metrics related to agentic AI security, aligned with the TRiSM pillars.
Attack Evaluation Benchmarks
Several benchmark datasets have emerged to evaluate adversarial attacks against LLM-based agents. For prompt injection and jailbreak scenarios, HarmBench [mazeika2024harmbench] and JailbreakBench [chao2024jailbreakbench] provide standardized datasets for evaluating robustness against adversarial prompts, including role-playing, obfuscation, and multi-turn manipulation scenarios particularly relevant to agentic systems. Tool misuse is addressed by ToolBench [xu2023tool] and API-Bank [li2023apibank], which evaluate agents’ ability to use external tools correctly while resisting adversarial instructions that trigger unauthorized access or malicious function calls.
Multi-agent adversarial scenarios are covered by AgentBench [liu2023agentbench] and GAIA [mialon2023gaia], which test resilience against coordinated attacks, impersonation, and information manipulation in multi-agent environments. Recent datasets also evaluate vulnerabilities specific to agentic architectures, including memory poisoning attacks where adversaries inject malicious information into agent memory stores, and context manipulation where attack payloads are embedded in retrieved documents or tool outputs [raza2025justhumansneedvaccines].
Defense Evaluation Metrics
Standard metrics for assessing defense mechanisms have been established across the research community. Attack Success Rate (ASR) measures the percentage of attack attempts that successfully compromise agent behavior, bypass safety constraints, or extract sensitive information, with lower ASR indicating more robust defenses [saha2025breaking]. Robustness Score quantifies the agent’s ability to maintain intended behavior under adversarial conditions, typically measured as the degradation in task performance when subjected to attacks compared to benign scenarios [raza2025humanibench].
Defense Effectiveness Rate captures the proportion of attacks successfully detected and mitigated by defense mechanisms, including input validation, output filtering, and behavioral monitoring systems. For production deployments, False Positive and False Negative Rates are critical, measuring the rate at which legitimate user requests are incorrectly flagged as attacks and actual attacks that evade detection, respectively. Our proposed Composite Security Score (CSS), introduced in Section 5 provides a holistic assessment by combining multiple security dimensions including input validation coverage, tool access control effectiveness, memory integrity, and inter-agent authentication strength.
Table 8 provides a systematic comparison of these benchmarks, including their coverage of attack scenarios, applicable defense metrics, and alignment with TRiSM pillars.
Table 8: Evaluation Benchmarks for Agentic AI Security
Benchmark Attack Coverage Metrics TRiSM Pillars Limitations
HarmBench [mazeika2024harmbench] Prompt injection, jailbreaks ASR, robustness Security, Explainability Limited multi-agent
ToolBench [xu2023tool] Tool misuse, unauthorized API Effectiveness, FP/FN Security, Privacy Tool-focused only
AgentBench [liu2023agentbench] Multi-step attacks, coordination Task success, compliance All pillars Limited adversarial
GAIA [mialon2023gaia] Ambiguous instructions, reasoning Completion, safety Explainability, Governance Early stage
WebArena [zhou2023webarena] Web attacks, malicious content Success, safety Security, Privacy Domain-specific
HELM [liang2022helm] Robustness, fairness, bias Comprehensive ModelOps, Governance Not agent-specific
MLCommons [mlcommons2024] Harmful content, privacy leakage Safety, privacy scores Security, Privacy, Governance Evolving
Key Insights
This section presented a defense-in-depth view of AMAS security spanning data confidentiality (encryption and secure enclaves), access control (RBAC/ABAC with least privilege and dynamic trust), adversarial resilience (prompt/tool-output validation and training-based hardening), and runtime monitoring (auditing, anomaly detection, and trust scoring). Privacy protection was framed as a system property: differential privacy to limit re-identification risk, data minimization through scoped/ephemeral memory and pseudonymization, and secure computation (SMPC, homomorphic encryption, and TEEs) to enable collaboration without raw-data exposure. We also summarized emerging evaluation practice, such as benchmarks for prompt injection, tool misuse, and multi-agent manipulation, together with metrics such as ASR, robustness degradation, defense effectiveness, and FP/FN rates; to quantify both vulnerabilities and mitigation quality in a TRiSM-aligned manner. Overall, security and privacy shift from optional add-ons to foundational requirements for regulated deployments.
7 Compliance with AI Governance and Policy Frameworks for Agentic AI Systems
@@ 7.1 Regulatory Standards
  • NIST AI Risk Management Framework (AI RMF) [tabassi2023ai]. A voluntary U.S. framework organized around the Govern, Map, Measure, Manage functions. It emphasizes organizational governance, risk identification and measurement, and operational risk treatments. Mapping to the RMF also benefits from NISTIR 8312 (Explainable AI principles) for transparency obligations.
  • EU AI Act [eu_ai_act]. A comprehensive EU regulation with phased application: bans on certain practices apply after six months (Feb 2025), obligations for general-purpose AI begin Aug 2025, and most high-risk system obligations (e.g., risk management, documentation, human oversight, post-market monitoring) start Aug 2026. High-risk systems require ongoing risk assessment, technical documentation (Annex IV), logging/traceability (Arts. 11–12), data governance (Art. 10), human oversight (Art. 14), and robustness/accuracy (Art. 15).
@@ 7.2 Auditability
AMAS can produce emergent and opaque behaviors; auditors must reconstruct the chain of decisions.
@@ 7.4 Documentation & Record-Keeping (Evidence)
  • Technical documentation (EU AI Act [EU_AI_Act2025]): system purpose and risk class, training/eval data governance, performance, limitations, foreseeable misuse, human oversight measures.
@@ 7.5 Post-Market Monitoring & Incident Response
  • Periodic reassessment. Re-run impact/risk assessments (ISO/IEC 42005 [ISO_IEC_42005_2025], ISO/IEC 23894 [ISOIEC23894_2023] ) when capabilities, data, or context change materially.
@@ 7.6 Supply-Chain & Secure Development (GenAI)
  • OWASP LLM risks. Mitigate prompt injection, insecure output handling, data poisoning, model DoS, and supply-chain risks using the OWASP [owasp2024llmtop10] Top 10 for LLM applications.
@@
Key Insights
This section operationalized compliance through systematic mappings between TRiSM pillars and regulatory instruments (EU AI Act, NIST AI RMF, ISO/IEC 42001, GDPR, HIPAA). Comprehensive auditability via immutable logging and decision provenance graphs enables regulatory audit and incident investigation. Policy-as-code mechanisms encode runtime constraints, preventing violations through pre-execution validation and human sign-off requirements for high-impact operations. Documentation rigor (technical dossiers, model cards, evaluation reports) provides regulator-ready evidence packages, while post-market monitoring implements continuous telemetry, incident handling workflows, and periodic reassessment protocols. Supply chain security addresses provenance verification through Software Bill of Materials documentation and dependency scanning. This integrated lifecycle program transforms compliance from checkbox exercise to strategic capability, building evidenceable trust through systematic controls and adaptive risk management.
@@ 8 Discussion
In this section, we summarize key findings and outline implications and future directions for deploying TRiSM in AMAS.
@@ 8.1 Technical Implications of TRiSM for Agentic AI Design
Rather than treating LLM agents as black-box decision-makers, TRiSM encourages designing agents with continuous oversight guardrails [guardrails_ai_guardrails_2024]. Recent discussions also highlight the value of specialized “guardian agents” within AMAS [securiti2025aitrism], which can filter sensitive data, establish baselines of normal behavior, and enforce runtime policies (e.g., blocking disallowed actions such as exposing personally identifiable information). Prior work also warns about “excessive agency” [owasp2025agentic], where overly autonomous agents with broad tool access may trigger unintended harmful actions (e.g., hallucination-driven execution or misinterpreted goals). TRiSM mitigates these risks by limiting autonomy through role separation, scoped permissions, and safety constraints. Emerging AMAS-specific threats such as prompt injection, memory poisoning, and cascading hallucinations further motivate built-in risk controls.
8.2 Ethical and Societal Ramifications
Beyond technical risks, deploying networks of autonomous agents raises ethical and societal concerns around accountability, fairness, and human oversight [raza2025responsible]. TRiSM governance emphasizes that organizations must retain responsibility for AI actions and avoid obscuring accountability behind opaque decision-making. In practice, each autonomous agentic decisions should be transparent enough to be understood and challenged by human reviewers when necessary [vayani2025all]. TRiSM does not remove humans from the loop; instead, it structures human-agent collaboration through defined oversight roles and escalation pathways [raza2025humanibench]. In AMAS deployments, humans may monitor agent swarms in real time, pause or shut down anomalous behavior, and adjust policies on the fly. The risk of user complacency, over-trusting autonomous agents; has also been noted as a hazard [parasuraman2010complacency], and TRiSM counteracts this by formalizing oversight and supervision requirements.
8.3 Critical Analysis of Evaluation Approaches
The evaluation framework in Sections 5 and 6.3 supports targeted diagnosis by separating trustworthiness, explainability, user-centered performance, and coordination metrics, rather than relying on monolithic accuracy [raza2024exploring]. The proposed CSS and TUE further address gaps in capturing multi-agent phenomena that single-agent metrics cannot represent [ribeiro2016should, lundberg2017unified]. However, several limitations remain. Human-centered evaluation is resource-intensive and does not scale to continuous deployment, and AMAS evaluation often lacks ground truth for emergent coordination behaviors. Improvements in one dimension may degrade another (e.g., verbose explanations increasing latency), and composite weights vary by domain, requiring clearer guidance for practitioners. Current protocols also struggle to capture behavioral drift over time, and subjective measures often omit inter-rater reliability. Finally, evaluation metrics may be gamed (Goodhart’s Law [^21] Goodhart’s law.), especially when systems optimize for a single composite score rather than underlying capability.
8.4 Implementation Challenges and Real-World Constraints
While TRiSM provides a strong foundation, deploying it in production introduces practical constraints. Continuous monitoring, validation, and anomaly detection add computational overhead and latency, which can harm user experience in time-sensitive settings. Large-scale enterprise deployments may involve hundreds or thousands of agents, increasing compute, storage, networking, and staffing costs. Integration is also challenging because many organizations rely on legacy infrastructure or third-party frameworks not designed for agentic security; adding memory protection, access control, and behavioral monitoring often requires cross-component changes and ongoing maintenance. Operationally, monitoring can produce large volumes of alerts, including false positives, which may overwhelm security teams. A phased, risk-driven adoption strategy is therefore often more feasible: start with input validation, restricted tool access, and logging, and then add stronger controls as system maturity and risk exposure increase.
@@ 8.5 Alignment with Emerging AI Regulations and Standards
TRiSM principles align closely with emerging regulatory frameworks and standards. For example, the EU AI Act [eu_ai_act] emphasizes risk management, transparency, data governance, and human oversight for high-risk systems, for example capabilities directly supported by TRiSM-aligned governance and controls [yang2025agentic, farooq2025evaluating]. This convergence suggests that TRiSM-based practices can help organizations translate technical safeguards into compliance-ready processes and evidence.
Despite progress, some gaps remain. For example, there is still no consensus on standardized benchmarks and metrics for evaluating agentic systems under TRiSM principles, which limits cross-study comparison and objective progress tracking. Many controls also lack real-world validation beyond lab settings or narrowly scoped tasks, and adversarial learning remains an ongoing challenge as attackers adapt to defenses. Future work should prioritize standardized stress-testing protocols (e.g., red-teaming), longitudinal evaluation to detect drift, and deployment case studies that report both effectiveness and operational costs.
Extension to Multimodal and Embodied Agents. While this survey focuses on LLM-based AMAS, the Agentic AI landscape increasingly encompasses multimodal agents (e.g., vision-language models such as LLaVA, GPT-4V) and embodied agents [xia2018gibson] operating in physical environments (e.g., robotic systems, autonomous vehicles, industrial automation). Future work should systematically extend TRiSM frameworks to address the unique challenges these systems present. Multimodal agents introduce additional security considerations including cross-modal adversarial attacks (e.g., adversarial images triggering unsafe text outputs), visual hallucinations, and sensor spoofing vulnerabilities. Embodied agents face physical safety risks, real-time decision-making constraints under uncertainty, sim-to-real transfer gaps, and complex human-robot trust calibration requirements. Extending the proposed CSS and TUE metrics to capture sensor fusion quality, physical action safety scores, and human-robot collaboration efficacy represents a promising direction for comprehensive agentic AI evaluation.
@@ 8.6 Future Roadmap for Agentic AI TRiSM
The community should create open benchmarks and challenge environments to test AMAS governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed ASAM that performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” [bacudio2011overview], and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [schneider2020wargaming] , can help discover vulnerabilities in a controlled way before real adversaries do.
We encourage designing better interfaces and protocols for human oversight of Agentic AI systems. Borrowing from human-computer interaction [AmershiWVFNCSIB19] and cognitive engineering [hexmoor2025behaviour], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
@@ 9 Conclusion
Agentic AI systems built as AMAS are rapidly changing how autonomous systems collaborate, plan, and make decisions in high-stakes domains such as healthcare, finance, and public services. At the same time, these systems expand the risk surface and raise new challenges in trust, risk, and security management (TRiSM). This review addressed those challenges and introduced a TRiSM-oriented framework for LLM-based AMAS organized around five pillars: Explainability, ModelOps, Security, Privacy, and lifecycle Governance. We contributed a risk taxonomy that captures AMAS-specific threats, including prompt injection, memory poisoning, and collusive or cascading failures. We also proposed practical evaluation metrics to measure inter-agent enablement and the correctness and efficiency of tool use. In addition, we summarized key technical mechanisms to support TRiSM in practice, including explainability methods and security/privacy protections. Looking ahead, progress will depend on stronger adversarial robustness, clearer governance and audit protocols, and standardized benchmarks that evaluate both trustworthiness and coordination under realistic deployment conditions. Aligning Agentic AI development with TRiSM principles can help ensure these systems remain not only capable and scalable, but also safe, transparent, and accountable for real-world use.
Acknowledgments
This research was funded by the European Union’s Horizon Europe research and innovation programme under the AIXPERT project (Grant Agreement No. 101214389), which aims to develop an agentic, multi-layered, GenAI-powered framework for creating explainable, accountable, and transparent AI systems. Additionally, this work was supported in part by the National Science Foundation (NSF); in part by United States Department of Agriculture (USDA); in part by the National Institute of Food and Agriculture (NIFA), through the “Artificial Intelligence (AI) Institute for Agriculture” Program, Accession Number 1029004 for the Project Titled “Robotic Blossom Thinning with Soft Manipulators” under Award AWD003473, Award AWD004595, and Award USDA-NIFA; and in part by United States Department of Agriculture National Science Foundation (USDANSF), Accession Number 1031712, under the Project “ExPanding University of Central Florida (UCF) AI Research To Novel Agricultural EngineeRing Applications (PARTNER)” under Grant 2024-67022-41788.